"The Length of a rectangular label is 5 centimeters less than 2 times the width. The perimeter is 110 centimeters. find the width."
I am having trouble solving this problem and am not sure where to begin.

Answer by rfer(16322)

2x+2(2x-5)=110
2x+4x-10=110
6x=120
x=120/6
x=20 cm width